# EfficientQAT: Efficient Quantization-Aware Training for Large Language Models

### v1

@@ 4.1 EfficientQAT for LLMs Quantization
Evaluation. We assess the zero-shot accuracy of five common-sense reasoning tasks using the v0.4.2 lm-evaluation-harnesshttps://github.com/EleutherAI/lm-evaluation-harness. The tasks include WinoGrande winogrande , PIQA piqa , HellaSwag hellaswag , Arc-Easy arc , and Arc-Challenge arc . We also measure the perplexity of Wikitext2 and C4 with a 2048 context length, as done in previous studies gptq ; omniquant .
PTQ Baseline. We compare our results with PTQ methods from uniform quantization such as GPTQ gptq , AWQ awq , OmniQ omniquant , and AutoRound weight_round , and vector quantization including QuIP# quip-sharp and AQLM aqlm .
@@
PTQ Results. The accuracy and perplexity results are presented in Table 1 and Table 3, respectively. It seems that minimal performance degradation with 4-bit group quantization, and even simple round-to-nearest (RTN) quantization achieves respectable results. Therefore, EfficientQAT offers a slight perplexity improvement ($\sim$0.02) in 4-bit uniform quantization and comparable zero-shot accuracy with existing methods. As the number of quantization bits decreases, EfficientQAT’s performance advantage increases. In 3-bit quantization, EfficientQAT surpasses the optimization-based method AutoRound by approximately 0.5% in zero-shot accuracy and outperforms OmniQuant by 0.14 to 0.43 in perplexity. Notably, in 2-bit quantization, EfficientQAT outperforms AutoRound weight_round by 5% in zero-shot accuracy for Llama-2-7b. EfficientQAT also exceeds previous vector quantization methods like AQLM aqlm using a 2x8 codebook. While AQLM aqlm with a 1x16 codebook and QuIP# quip-sharp perform slightly better in 2-bit scenarios, they introduce notable computational overhead llm-qbench and can even reduce inference speed, as shown in Figure 2(a). Thus, EfficientQAT successfully pushes the limits of uniform quantization in LLMs, making hardware-friendly uniform quantization competitive with more complex vector quantization methods. However, we can also find that Llama-3 models suffer more performance degeneration after quantization llama3_quant , which may caused by the full training with 15T tokens other than 2T tokens.
QAT Baseline. We also compare our results with existing QAT methods, including LLM-QAT llmqat , BitDistiller bitdistiller , PB-LLM pbllm and DB-LLM dbllm .
QAT Results. As demonstrated in Table 2, EfficientQAT markedly surpasses other QAT methods in performance. Specifically, EfficientQAT reduces perplexity by 0.89 compared to BitDistiller bitdistiller when applied to the Llama-2-7B model with w2g128 quantization. In the more challenging scenario of Llama-3 quantization llama3_quant , EfficientQAT exceeds the performance of DB-LLM dbllm by 5.98 in C4 perplexity and by 8.57 points in average accuracy with the w2g128 quantization.
@@
OmniQ 2 uniform 128 11.06 8.26 6.11 - - 15.02 11.05 8.52 - -
@@ 4.2 EfficientQAT for Instruction Tuning
Training and Evaluation. Following existing works qa-lora ; ir-qlora , we train Llama-1 models on the Alpaca dataset alpaca and assess their performance by measuring average 5-shot MMLU mmlu accuracy works qa-lora ; ir-qlora . The training hyperparameters are identical to those described in Section 4.1, except we replace the RedPajama dataset redpajama with Alpaca. In line with QLoRA’s methodology qlora , we adjust the source context length to 384 and the target context length to 128, training for 10,000 steps with a batch size of 16.
Baseline. We benchmark EfficientQAT against several leading methods, including QLoRA qlora , QA-LoRA qa-lora , PEQA peqa , and IR-QLoRA ir-qlora , across quantization setting of 2, 3, and 4 bits. Consistent with QA-LoRA qa-lora , we also employ GPTQ gptq to quantize the fine-tuned QLoRA models into a low-bit format without FP16 LoRA for equitable comparison.
Results. Both Table 7 and Figure 1(b) indicate that EfficientQAT significantly outperforms existing Q-PEFT methods. For instance, in channel-wise quantization (group size of -1), EfficientQAT achieves more than 3% higher accuracy than PEQA peqa . In the 2-bit quantization scenario, the superiority of EfficientQAT is even more pronounced, surpassing QA-LoRA qa-lora by 5.1% and 4.0% in 7B and 13B models, respectively, and outperforming PEQA by 4.5% and 8.7% in the same models. Moreover, Table 7 also demonstrates that EfficientQAT outperforms both QA-LoRA and QLoRA with GPTQ in smaller model memory footprint (larger group size).
@@ 4.3 Ablation Analysis
Effectiveness of each component. As indicated in Table 7, both the Block-AP and E2E-QP components significantly enhance performance, with their combination yielding the best results. Notably, Block-AP outperforms E2E-QP, aligning with findings from BRECQ brecq .
Trainable parameters of Block-AP. Block-AP trains all parameters, including original weights and quantization parameters. Previous methods have introduced various training strategies to mitigate overfitting, such as trained rounding adaround ; weight_round , clipping thresholds omniquant , and step sizes lsq ; cbq . We compare Block-AP with these methods by modifying only the trainable parameters of Block-AP. As shown in Table 7, Block-AP (training $s$, $z$, $\mathbf{W}$) performs best with an acceptable training cost. Additionally, the memory footprint of directly training $\mathbf{W}$ is even smaller than that of training the rounding operation, which requires an additional copy of rounding parameters.
@@
Table 9: Speed of the FP16 linear layer matrix-vector multiplication in PyTorch, and relative INT2 speedups in BitBLAS bitblas . Testing on A100-80GB GPU.
@@
4.4 Efficiency of EfficientQAT
Training Efficiency Table 8 illustrates the required memory and time for training Lllama-2 models using EfficientQAT. The results indicate that the model completes training rapidly, taking 4.8 hours for the 7B model and 40.9 hours for the 70B model. Furthermore, the adoption of a quantized backbone in the end-to-end quantization process (E2E-QP) significantly reduces memory usage due to the lower bit width. Consequently, EfficientQAT accomplishes 2-bit training for the 70B model using only 34.2GB of memory.
Inference Efficiency Weight-only quantization reduces the memory requirements of LLMs and accelerates inference, as LLMs are predominantly memory-bound llm-unveil . Quantization modifies the computation in the linear layers while other operations remain unchanged. To evaluate this, we calculate the relative speedup in the linear layers’ processing. According to Table 9, INT2 quantization enhances the forward-pass speed by approximately 2.9x to 4.4x. Furthermore, due to the leverage of standard uniform quantization, the quantized models of EfficientQAT can also achieve speedup through other toolbox, such as MLC-LLM mlc-llm , AWQ awq , and BitBLAS bitblas .
5 Conclusion
Appendix A Limitations
In terms of limitations, EfficientQAT demonstrates marked superiority in low-bit scenarios, such as 3-bit and 2-bit quantization. However, in the 4-bit group-wise quantization, existing PTQ methods like GPTQ [22], AWQ [37], and even the simplest RTN achieve comparable performance more quickly. Additionally, the Llama-3 quantization experiences significant performance degradation [28], a phenomenon also observed in EfficientQAT. Moreover, the 2-bit performance of EfficientQAT still exhibits a small gap (3% to 4% accuracy) compared to full-precision models. Addressing these issues to achieve nearly lossless performance with INT2 quantization is a key focus of our future work.
Appendix B Broader Impact
This paper presents work whose goal is to advance the compression and acceleration of large language models. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix C Gradient of Trainable Parameters in Block-AP
Block-AP, aligned with LSQ+[4], uses a straight-through estimator (STE)[3] to facilitate gradient computation through the rounding operation. The gradients of scaling factor $s$ are computed as follows:
and the gradient with respect to zero point $z$ is:
[…]
@@ Appendix C Gradient of Trainable Parameters in Block-AP
In this study, we present a thorough comparison of our method against existing PTQ techniques, including GPTQ [22], AWQ [37], OmniQ [52], AutoRound [11], QuIP# [58], and AQLM [20]. We also compare with existing QAT methods, including LLM-QAT [43], BitDistiller [19], PB-LLM [51] and DB-LLM [9]. Additionally, we also evaluate quantized parameter-efficient fine-tuning methods such as PEQA [29], QLoRA [16], QA-LoRA [64], and IR-QLoRA [47]. The results we discuss originate from their respective official publications, and other scholarly articles, or are derived from our reproduction. We meticulously document the source of the results for each method as follows:
  • GPTQ, AWQ, OmniQ, AutoRound: The zero-shot accuracy results for Llama-2 models using these methods are derived from the AutoRound GitHub repository§§§AutoRound: https://github.com/intel/auto-round/blob/main/docs/acc.md. The perplexity results for the Llama-2 models using GPTQ, AWQ, and OmniQ are taken from the OmniQ paper [52]. The results for Llama-3 models using AWQAWQ:https://github.com/mit-han-lab/llm-awq and GPTQGPTQ:https://github.com/qwopqwop200/GPTQ-for-LLaMa were obtained through their open-source implementations.
  • QuIP#, AQLM: We replicated the results using the official pre-trained models provided by QuIP#******https://github.com/Cornell-RelaxML/quip-sharp and AQLM††††††https://github.com/Vahe1994/AQLM.
  • LLM-QAT, BitDistiller: These results are cited from BitDistiller [19] paper.
  • PB-LLM, DB-LLM: These results are cited from recent Llama-3 quantization empirical study [28].
  • PEQA: The per-channel quantization results (g=-1) are cited from their publication [29], and the results for a group size of 64 were produced using our codebase.
  • QA-LoRA, QLoRA, QLoRA w/ GPTQ: These results are cited from QA-LoRA [64] paper.
  • IR-QLoRA: These results are cited from IR-QLoRA [47] paper.
@@ Appendix F Additional Ablation Analysis
Table 12: Results about instruction tuning of large vision-language models. We following the overall training pipeling of LLaVA-1.5 [39] and just change the fine-tuning methods. ‘QLoRA + Block-AP’ indicates that we leverage proposed Block-AP to quantized the QLoRA models into low-bits for fair comparisons. MME’s perception scores are normalized to 100 percent.
@@
Traditional Q-PEFT methods only do experiments on the language models. In this section, we further extend proposed EfficientQAT into Large vision-Language models (LVLMs) such as LLaVA [40].
Training and Evaluation. For the fine-tuning of large vision-language models (LVLMs), we largely align with LLaVA1.5 [39], which encompass the training model, datasets, and hyperparameters‡‡‡‡‡‡For comprehensive details, please consult the official repository at https://github.com/haotian-liu/LLaVA.. Unlike LLaVA1.5, which begins fine-tuning with full-precision Vicuna models using either full fine-tuning or LoRA-based methods [27], EfficientQAT starts with Vicuna models already quantized using our Block-AP method and continues with our E2E-QP fine-tuning approach. The training process involves two steps: initially freezing the LLM and pre-training a projector to align features with a Vision Transformer (ViT), followed by end-to-end fine-tuning of both the LLM and the projector. For EfficientQAT, we modify the learning rates in the second step to $2\times 10^{-5}$ for 4-bit and $3\times 10^{-5}$ for 2-bit and 3-bit.
Evaluation. Evaluation of the fine-tuned LVLMs are conducted across four benchmarks: MME [23], MM-Vet [67], MMBench [42], and ScienceQA [44].
Baseline. We compare our results with those of QLoRA [16], applying our Block-AP method to quantize the QLoRA fine-tuned models to low bits for fair comparison.
Results. As shown in Table 12, EfficientQAT outperforms QLoRA [16] in low-bit settings for both LLaVA-1.5-7B and LLaVA-1.5-13B models, consistent with previous results in LMMs. Remarkably, the 2-bit LLaVA-1.5-13B model trained with EfficientQAT achieves an average score of 59.9, surpassing the 59.6 of the FP16 LLaVA-1.5-7B model trained with LoRA. However, there is a slight performance decrease observed in the 4-bit EfficientQAT and 16-bit QLoRA compared to the 16-bit LoRA, indicating that further research is needed to optimize Q-PEFT within LVLMs.

### v2

@@ 4.1 EfficientQAT for LLMs Quantization
3-70B PB-LLM 2 128 11.60 34.50 46.00
3-70B EfficientQAT 2 128 6.07 9.23 67.57
Evaluation. We assess the zero-shot accuracy of five common-sense reasoning tasks using the v0.4.2 lm-evaluation-harness***https://github.com/EleutherAI/lm-evaluation-harness. The tasks include WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), Arc-Easy (Clark et al., 2018), and Arc-Challenge (Clark et al., 2018). We also measure the perplexity of Wikitext2 and C4 with a 2048 context length, as done in previous studies (Frantar et al., 2022; Shao et al., 2023).
PTQ Baseline. We compare our results with PTQ methods from uniform quantization such as GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2023), OmniQ (Shao et al., 2023), ApiQ (Liao & Monz, 2024) and AutoRound (Cheng et al., 2023), and vector quantization including QuIP# (Tseng et al., 2024) and AQLM (Egiazarian et al., 2024). Note that if a result is the best of uniform quantization, we set it to bold.
PTQ Results. The accuracy and perplexity results are presented in Table 1 and Table 3, respectively. It seems that minimal performance degradation with 4-bit group quantization, and even simple round-to-nearest (RTN) quantization achieves respectable results. Therefore, EfficientQAT offers a slight perplexity improvement ($\sim$0.02) in 4-bit uniform quantization and comparable zero-shot accuracy with existing methods. As the number of quantization bits decreases, EfficientQAT’s performance advantage increases. In 3-bit quantization, EfficientQAT surpasses the optimization-based method AutoRound by approximately 0.5% in zero-shot accuracy and outperforms OmniQuant by 0.14 to 0.43 in perplexity. Notably, in 2-bit quantization, EfficientQAT outperforms AutoRound (Cheng et al., 2023) by 5% in zero-shot accuracy for Llama-2-7b. EfficientQAT also exceeds previous vector quantization methods like AQLM (Egiazarian et al., 2024) using a 2x8 codebook. While AQLM (Egiazarian et al., 2024) with a 1x16 codebook and QuIP# (Tseng et al., 2024) perform slightly better in 2-bit scenarios, they introduce notable computational overhead (Gong et al., 2024) and can even reduce inference speed. Thus, EfficientQAT successfully pushes the limits of uniform quantization in LLMs, making hardware-friendly uniform quantization competitive with more complex vector quantization methods. However, we can also find that Llama-3 models suffer more performance degeneration after quantization (Huang et al., 2024), which may caused by the full training with 15T tokens other than 2T tokens.
QAT Baseline. We also compare our results with existing QAT methods, including LLM-QAT (Liu et al., 2023e), BitDistiller (Du et al., 2024), PB-LLM (Shang et al., 2023) and DB-LLM (Chen et al., 2024).
QAT Results. As demonstrated in Table 2, EfficientQAT markedly surpasses other QAT methods in performance. Specifically, EfficientQAT reduces perplexity by 0.89 compared to BitDistiller (Du et al., 2024) when applied to the Llama-2-7B model with w2g128 quantization. In the more challenging scenario of Llama-3 quantization (Huang et al., 2024), EfficientQAT exceeds the performance of DB-LLM (Chen et al., 2024) by 5.98 in C4 perplexity and by 8.57 points in average accuracy with the w2g128 quantization.
@@
OmniQ 2 uniform 128 11.06 8.26 6.55 - - 15.02 11.05 8.52 - -
ApiQ 2 uniform 128 8.25 6.71 - - - 12.04 9.13 - - -
@@ 4.2 EfficientQAT for Instruction Tuning
Training and Evaluation. Following existing works (Xu et al., 2023b; Qin et al., 2024), we train Llama-1 models on the Alpaca dataset (Taori et al., 2023) and assess their performance by measuring average 5-shot MMLU (Hendrycks et al., 2020) accuracy works (Xu et al., 2023b; Qin et al., 2024). The training hyperparameters are identical to those described in Section 4.1, except we replace the RedPajama dataset (Computer, 2023) with Alpaca. In line with QLoRA’s methodology (Dettmers et al., 2023a), we adjust the source context length to 384 and the target context length to 128, training for 10,000 steps with a batch size of 16.
Baseline. We benchmark EfficientQAT against several leading methods, including QLoRA (Dettmers et al., 2023a), QA-LoRA (Xu et al., 2023b), PEQA (Kim et al., 2023a), and IR-QLoRA (Qin et al., 2024), across quantization setting of 2, 3, and 4 bits. Consistent with QA-LoRA (Xu et al., 2023b), we also employ GPTQ (Frantar et al., 2022) to quantize the fine-tuned QLoRA models into a low-bit format without FP16 LoRA for equitable comparison.
Results. Both Table 7 and Figure 1(b) indicate that EfficientQAT significantly outperforms existing Q-PEFT methods. For instance, in channel-wise quantization (group size of -1), EfficientQAT achieves more than 3% higher accuracy than PEQA (Kim et al., 2023a). In the 2-bit quantization scenario, the superiority of EfficientQAT is even more pronounced, surpassing QA-LoRA (Xu et al., 2023b) by 5.1% and 4.0% in 7B and 13B models, respectively, and outperforming PEQA by 4.5% and 8.7% in the same models. Moreover, Table 7 also demonstrates that EfficientQAT outperforms both QA-LoRA and QLoRA with GPTQ in smaller model memory footprint (larger group size).
@@ 4.3 Ablation Analysis
Effectiveness of each component. As indicated in Table 7, both the Block-AP and E2E-QP components significantly enhance performance, with their combination yielding the best results. Notably, Block-AP outperforms E2E-QP, aligning with findings from BRECQ (Li et al., 2021).
Trainable parameters of Block-AP. Block-AP trains all parameters, including original weights and quantization parameters. Previous methods have introduced various training strategies to mitigate overfitting, such as trained rounding (Nagel et al., 2020; Cheng et al., 2023), clipping thresholds (Shao et al., 2023), and step sizes (Esser et al., 2019; Ding et al., 2023). We compare Block-AP with these methods by modifying only the trainable parameters of Block-AP. As shown in Table 7, Block-AP (training $s$, $z$, $\mathbf{W}$) performs best with an acceptable training cost. Additionally, the memory footprint of directly training $\mathbf{W}$ is even smaller than that of training the rounding operation, which requires an additional copy of rounding parameters.
@@
Table 9: Comparisons of training time with existing methods in Llama-2-70B.
Method One A100-80GB? GPU hours (h)
LLM-QAT 900
QuiP# 300
AQLM 336
BitDistiller 64
PB-LLM 90
DB-LLM 82
EfficientQAT 41
4.4 Efficiency of EfficientQAT
Training Efficiency Table 8 illustrates the required memory and time for training Lllama-2 models using EfficientQAT. The results indicate that the model completes training rapidly, taking 4.8 hours for the 7B model and 40.9 hours for the 70B model. we further compare the training time with other QAT methods, including BitDistiller, PB-LLM, and DN-LLM. As shown in Table 1(c), the training time of EfficientQAT is significantly lower than that of existing methods. For example, the tuning time of EfficientQAT is only 14% of AQLM and 50% of DB-LLM. Additionally, for quantizing a 70B model, the full process of EfficientQAT can be completed on a single A100-80GB GPU. However, other methods, except AQLM, require at least 4 A100-80GB GPUs to quantize a model of this size. Therefore, EfficientQAT is both a time-efficient and memory-efficient QAT method.
Inference Efficiency Due to the leverage of standard uniform quantization, the quantized models of EfficientQAT can also achieve speedup through a lot of toolboxes, such as MLC-LLM (team, 2023), AWQ (Lin et al., 2023), and BitBLAS (Wang et al., 2024), T-MAC (Wei et al., 2024), Marlin (Frantar et al., 2024), etc. For example, Table 10 shows that INT2 quantization of EfficientQAT can enhance the forward-pass speed by approximately 2.9x to 4.4x through BitBLAS (Wang et al., 2024).
5 Conclusion
Appendix A Reproducibility Statement
In this section, we summarize the necessary information to reproduce our results. We provide the training and evaluation details at the beginning of each sub-section in Sec. 4. We also provide the source of detailed results for each compared method in Sec.D.
Appendix B Gradient of Trainable Parameters in Block-AP
Block-AP, aligned with LSQ+(Bhalgat et al., 2020), uses a straight-through estimator (STE)(Bengio et al., 2013) to facilitate gradient computation through the rounding operation. The gradients of scaling factor $s$ are computed as follows:
and the gradient with respect to zero point $z$ is:
[…]
Appendix C Speedup with BitBlas
According to Table 10, INT2 quantization enhances the forward-pass speed by approximately 2.9x to 4.4x.
Table 10: Speed of the FP16 linear layer matrix-vector multiplication in PyTorch, and relative INT2 speedups in BitBLAS Wang et al. (2024). Testing on A100-80GB GPU.
@@ Appendix C Speedup with BitBlas
In this study, we present a thorough comparison of our method against existing PTQ techniques, including GPTQ (Frantar et al., 2022), AWQ (Lin et al., 2023), OmniQ (Shao et al., 2023), AutoRound (Cheng et al., 2023), QuIP# (Tseng et al., 2024), and AQLM (Egiazarian et al., 2024). We also compare with existing QAT methods, including LLM-QAT (Liu et al., 2023e), BitDistiller (Du et al., 2024), PB-LLM (Shang et al., 2023) and DB-LLM (Chen et al., 2024). Additionally, we also evaluate quantized parameter-efficient fine-tuning methods such as PEQA (Kim et al., 2023a), QLoRA (Dettmers et al., 2023a), QA-LoRA (Xu et al., 2023b), and IR-QLoRA (Qin et al., 2024). The results we discuss originate from their respective official publications, and other scholarly articles, or are derived from our reproduction. We meticulously document the source of the results for each method as follows:
  • GPTQ, AWQ, OmniQ, AutoRound: The zero-shot accuracy results for Llama-2 models using these methods are derived from the AutoRound GitHub repositoryAutoRound: https://github.com/intel/auto-round/blob/main/docs/acc.md. The perplexity results for the Llama-2 models using GPTQ, AWQ, and OmniQ are taken from the OmniQ paper (Shao et al., 2023). The results for Llama-3 models using AWQ§§§AWQ:https://github.com/mit-han-lab/llm-awq and GPTQGPTQ:https://github.com/qwopqwop200/GPTQ-for-LLaMa were obtained through their open-source implementations.
  • QuIP#, AQLM: We replicated the results using the official pre-trained models provided by QuIP#https://github.com/Cornell-RelaxML/quip-sharp and AQLM******https://github.com/Vahe1994/AQLM.
  • LLM-QAT, BitDistiller: These results are cited from BitDistiller (Du et al., 2024) paper.
  • PB-LLM, DB-LLM: These results are cited from recent Llama-3 quantization empirical study (Huang et al., 2024).
  • ApiQ: These results are cited from IR-ApiQ (Liao & Monz, 2024) paper.
  • PEQA: The per-channel quantization results (g=-1) are cited from their publication (Kim et al., 2023a), and the results for a group size of 64 were produced using our codebase.
  • QA-LoRA, QLoRA, QLoRA w/ GPTQ: These results are cited from QA-LoRA (Xu et al., 2023b) paper.
  • IR-QLoRA: These results are cited from IR-QLoRA (Qin et al., 2024) paper.
@@ Appendix E Size of Quantized Models
Table 13: Block-AP (w/o E2E-QP) results of Llama-2-7B in different calibration datasets.
Bits Dataset Wiki PPL C4 PPL Avg. Accuracy
w3g128 WikiText2 5.72 7.52 63.24
w3g128 C4 5.92 7.38 63.82
w3g128 Redpajama 5.91 7.41 63.50
[…]
w2g64 C4 7.87 9.30 59.24
w2g64 Redpajama 7.70 9.36 58.99
@@ Appendix F Additional Ablation Analysis
Training Dataset. More trainable parameters can increase the risk of overfitting. Previous works (Gong et al., 2024) show that a similar distribution between the calibration dataset and the test dataset can improve test accuracy. RedPajama and C4 datasets are diverse, while WikiText2 is simpler and sourced from Wikipedia. The close distribution of training and test datasets for WikiText2 results in significantly lower WikiText2 perplexity when using it as a calibration dataset. However, the average accuracy of zero-shot tasks in Table R7 shows that Block-AP’s generation ability is excellent, with only 0.26% and 1.28% accuracy declines when changing the calibration dataset from RedPajama to WikiText2 for w3g128 and w2g64, respectively. Additionally, using C4 as a calibration dataset can even increase the average accuracy by 0.2-0.3 points. Overall, we recommend using Block-AP with more diverse calibration datasets like C4 or RedPajama.
Table 14: Results about instruction tuning of large vision-language models. We following the overall training pipeling of LLaVA-1.5 Liu et al. (2023a) and just change the fine-tuning methods. ‘QLoRA + Block-AP’ indicates that we leverage proposed Block-AP to quantized the QLoRA models into low-bits for fair comparisons. MME’s perception scores are normalized to 100 percent.
@@
Traditional Q-PEFT methods only do experiments on the language models. In this section, we further extend proposed EfficientQAT into Large vision-Language models (LVLMs) such as LLaVA (Liu et al., 2023b).
Training and Evaluation. For the fine-tuning of large vision-language models (LVLMs), we largely align with LLaVA1.5 (Liu et al., 2023a), which encompass the training model, datasets, and hyperparameters††††††For comprehensive details, please consult the official repository at https://github.com/haotian-liu/LLaVA.. Unlike LLaVA1.5, which begins fine-tuning with full-precision Vicuna models using either full fine-tuning or LoRA-based methods (Hu et al., 2021), EfficientQAT starts with Vicuna models already quantized using our Block-AP method and continues with our E2E-QP fine-tuning approach. The training process involves two steps: initially freezing the LLM and pre-training a projector to align features with a Vision Transformer (ViT), followed by end-to-end fine-tuning of both the LLM and the projector. For EfficientQAT, we modify the learning rates in the second step to $2\times 10^{-5}$ for 4-bit and $3\times 10^{-5}$ for 2-bit and 3-bit.
Evaluation. Evaluation of the fine-tuned LVLMs are conducted across four benchmarks: MME (Fu et al., 2023), MM-Vet (Yu et al., 2023), MMBench (Liu et al., 2023d), and ScienceQA (Lu et al., 2022).
Baseline. We compare our results with those of QLoRA (Dettmers et al., 2023a), applying our Block-AP method to quantize the QLoRA fine-tuned models to low bits for fair comparison.
Results. As shown in Table 14, EfficientQAT outperforms QLoRA (Dettmers et al., 2023a) in low-bit settings for both LLaVA-1.5-7B and LLaVA-1.5-13B models, consistent with previous results in LMMs. Remarkably, the 2-bit LLaVA-1.5-13B model trained with EfficientQAT achieves an average score of 59.9, surpassing the 59.6 of the FP16 LLaVA-1.5-7B model trained with LoRA. However, there is a slight performance decrease observed in the 4-bit EfficientQAT and 16-bit QLoRA compared to the 16-bit LoRA, indicating that further research is needed to optimize Q-PEFT within LVLMs.